y=mx and 5=my if they are perpendicular find (m)
1
Expert's answer
2013-01-21T07:49:04-0500
If two lines are perpendicular then k1*k2=-1
We have y=mx and y=5/m
So in our case k1=m, k2=0
m*0=-1 is false for any value of m. Thus, there's no solution.